OpenStudy (anonymous):

(sinx+cosx)^2=1+2sinxcosx verify identity

OpenStudy (anonymous):

just expand (sinx + cosx)^2 = sin^2x + cos^2x + 2cosx sinx but sin^2x + cos^2x = 1. so..